# Broken detailed balance and entropy production in directed networks

### v4

@@ Abstract
The structure of a complex network plays a crucial role in determining its dynamical properties. In this work, we show that the the degree to which a network is directed and hierarchically organised is closely associated with the degree to which its dynamics break detailed balance and produce entropy. We consider a range of dynamical processes and show how different directed network features affect their entropy production rate. We begin with an analytical treatment of a 2-node network followed by numerical simulations of synthetic networks using the preferential attachment and Erdös-Renyi algorithms. Next, we analyse a collection of 97 empirical networks to determine the effect of complex real-world topologies. Finally, we present a simple method for inferring broken detailed balance and directed network structure from multivariate time-series and apply our method to identify non-equilibrium dynamics and hierarchical organisation in both human neuroimaging and financial time-series. Overall, our results shed light on the consequences of directed network structure on non-equilibrium dynamics and highlight the importance and ubiquity of hierarchical organisation and non-equilibrium dynamics in real-world systems.
@@ I Introduction
As anticipated, symmetry-breaking in the organisation of complex systems is the structural feature that drives network dynamical processes out of equilibrium. Dynamical processes on directed networks differ drastically from their undirected counterparts [38, 39] including in their phase-transitions [40], synchronisation properties [41, 42], topological resilience [43, 44] and pattern formation [45, 46]. Recent studies on extensive datasets of real-world networks, have shown a ubiquity of strong directedness and clear signs of hierarchical organisation [47, 48, 39, 49, 50]. As a result, the dynamics of directed networks are more indicative of the dynamics of real-world complex systems. In particular, this strong directedness results in a marked non-normality of the operators defined on such networks [51, 48]. Consequently, non-normality can result in the underlying networked systems diverging from equilibria [48]. Motivated by the ubiquity of directed structures in the real world and their significance in nonlinear dynamics, this paper aims to further strengthen this link using the formalism of stochastic thermodynamics.
@@ II Measures of directedness in networks
As anticipated, non-reciprocal interactions in complex systems disrupt detailed balance, leading to an increase in the EPR. This section delineates four established metrics to quantify the overall directedness of networks and to examine their influence on EPR within both random graphs and empirical datasets. These systems are represented as directed networks with non-negative weighted adjacency matrices, $\bm{W}=(W_{ij})$, where $W_{ij}\geq 0$ signifies the strength of the directed link $i\rightarrow j$, including self-edges $W_{ii}\geq 0$ and characterized by inherent asymmetry $\bm{W}\neq\bm{W}^{\top}$. By exploring four distinct metrics for network directedness, we seek to elucidate its effect on EPR, highlighting the emergence of hierarchical structuring within networks - an outcome deeply influenced by non-reciprocal interactions [62, 50].
@@ II.4 Parameterising directedness in networks
To investigate the dynamical affects of directedness on network dynamics, it would be preferable to continuously vary the level of asymmetry in a network. To this aim, we first generate highly hierarchical, non-normal networks using a preferential attachment (PA) scheme with weak reciprocal links (see Appendix B for details) [48]. We then linearly interpolate between this strongly directed network and its undirected Hermitian,
@@ III.1 Random walks on directed networks
which can, alternatively, be written as the the Kullback-Leibler (KL) divergence between the forward and backward joint transition rates between all pairs of states,
@@ III.2 The network-based multivariate Ornstein-Uhlenbeck process
$\frac{d\bm{x}}{dt}$ $=\Theta(\gamma\bm{W}-\bm{I})\bm{x}(t)+\bm{\nu}(t),$ (22)
@@ III.3 The Ising model
$P(\bm{x}(t+1)|\bm{x}(t))$ $=\prod_{i}\frac{\exp\left(x_{i}(t+1)\sum_{j}W_{ij}x_{j}(t)/T\right)}{2\cosh\left(\sum_{j}W_{ij}x_{j}(t)/T\right)},$ (32)
@@ III.3.2 Large networks of Ising spins
The combinatorial explosion of state space with the number of spins, means that estimating the steady state probabilities is computationally challenging for large systems. Alternative approaches include coarse-graining the state space by clustering together states and estimating transition probabilities [85, 14, 23], or using mean-field approximations [83]. Whilst using a coarse-grained state-space has been shown to capture the affect on the EPR of changing the temperature of the Ising model [23], it proved to be an inaccurate when varying the directedness of the network (see Appendix G) and so we hereby opt for a mean-field approach.
@@
$m_{i}(t)$ $\approx\tanh\sum_{j}W_{ij}m_{j}(t-1),$ (40)
$D_{il}(t)$ $\approx W_{il}(1-m^{2}_{i}(t))(1-m^{2}_{l}(t-1)).$ (41)
@@ IV.1 An exactly treatable case: 2-node networks
Finally, we can consider the Ising dynamics on this network. Whilst we do not have the exact solution, there are only four possible spin configurations in this system, therefore we can accurately estimate the steady state probabilities and consequently the EPR. The EPR, $\Phi$, for both the OU and Ising systems is plotted as a function of the network weights $(W_{12},W_{21})$ as shown in Figure 3. The positive diagonal $W_{12}=W_{21}$ represents the subset of undirected networks, which preserve detailed balance. In both cases, as the asymmetry of the network increases the EPR increases. The EPR in the OU is a function purely of the distance between the two weights, whereas for the Ising model, the relationship appears to be nonlinear.
@@ IV.2 Entropy production in parameterised synthetic networks
Equipped with a range of stochastic network dynamics where we can estimate the EPR, as well as a mechanism to systematically vary the directedness of the network, we can investigate the affect of directedness on the degree of non-equilibrium. Considering networks of sizes $N=$ 10, 100, 500 and 1000, we generate hierarchical networks using preferential attachment (see Appendix B) and then interpolate between these networks and their Hermitian as described in Section II. Recall the interpolation is given by,
@@ VII Entropy production and directed networks from multivariate time-series
The key limitation of such an approach is the assumption of the linear model for potentially nonlinear data. However, previous studies have found that highly nonlinear systems, such as large-scale brain dynamics [107], or even chaotic systems [108], are well approximated by linear models. Furthermore, autoregressive models have previously found success in a spectrum of areas including finance and economics [57], neuroscience [109] and beyond. Finally, we note that three recent studies in neuroimaging have considered related approaches, fitting linear models to neural recordings in order to quantify the EPR [97] or the asymmetry of interactions [98, 59] but that inferring time-reversibility from incomplete linear data may yield inaccuracies [96].
@@ VII.1 Linear auto-regression of multivariate time-series
$\bm{x}(t_{i+1})$ $=[\bm{I}+\Delta t(\bm{W-I})]\bm{x}(t_{i})+\bm{\Gamma}\bm{\zeta}_{i},$ (62)
@@ Appendix C Linear interpolation parameterises trophic directedness and irreciprocity
For original network, $\bm{h}$ is the vector of trophic levels which solves,
@@ Appendix D Deriving the entropy production rate in the Ornstein-Uhlenbeck process
$\bm{B}$ $=\begin{pmatrix}\Theta&\Theta(1-\gamma\tilde{W}_{12}(\epsilon))\\ \Theta(1-\gamma\tilde{W}_{21}(\epsilon))&\Theta\end{pmatrix},$ (132)
@@ Appendix F Circulant networks
It is important to note that $\bm{W}$ is circulant and therefore normal. However, as it is directed, it breaks detailed balance for all the systems considered in this paper (with the exception of the RW for $N=2$). More importantly, as we will show, applying the interpolation to the cyclic networks, $\epsilon$ increases the EPR of the corresponding OU. We found in Sections V and VI that the non-normality of the underlying network was a extremely strong indicator of the EPR of the NOU, yet in this case, the non-normality is 0 and the EPR still varies with $\epsilon$. This represents an important special case where hierarchical asymmetry or a global direction are not necessary for the OU to progressively break detailed balance. The lack of these cyclic structures in real-world networks [88] explains why, despite this special case, the non-normality correlates so closely with the EPR for the OU on real-world networks, as shown in Section VI.
@@
Therefore, the friction matrix $\bm{B}=\Theta(\bm{I}-\gamma\bm{\hat{W}}(\epsilon))$ is also circulant and defined by the vector,
$\bm{b}(\epsilon)=(\Theta,-\Theta\gamma\tilde{w}_{1}(\epsilon),\ldots,-\Theta\gamma\tilde{w}_{N-1}(\epsilon)).$ (143)
@@
$\tilde{b}_{k}$ $=\Theta-\Theta\gamma\sum_{j=1}^{N-1}\tilde{w}_{j}(\epsilon)\exp{\bm{i}\frac{2\pi jk}{N}}.$ (145)
@@
$\tilde{b}_{k}$ $=\Theta-\Theta\gamma\sum_{j=1}^{N-1}(\epsilon w_{j})(\bm{i}\sin(\frac{2\pi jk}{N})+\cos(\frac{2\pi jk}{N}))-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(1-\epsilon)(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N}),$ (148)
@@
$\Phi(\epsilon)$ $=\sum_{k=0}^{N-1}\frac{\Theta^{2}\gamma^{2}\left(\sum_{j=1}^{N}\epsilon w_{j}\sin(\frac{2\pi jk}{N})\right)^{2}}{\Theta-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(1-\epsilon)(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N})-\Theta\gamma\sum_{j=1}^{N-1}\epsilon w_{j}\cos(\frac{2\pi jk}{N})},$ (149)
$=\epsilon^{2}\sum_{k=0}^{N-1}\frac{\Theta^{2}\gamma^{2}\left(\sum_{j=1}^{N}w_{j}\sin(\frac{2\pi jk}{N})\right)^{2}}{\Theta-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N})},$ (150)
@@ Appendix H Derivation of the naive mean field for the asymmetric Ising model
$h^{\alpha}_{i}(t+1)=(1-$ $\alpha)\Theta_{i}(t)+\alpha\left(\frac{1}{T}(H_{i}+\sum_{j}W_{ij}x_{j}(t))\right).$ (159)
@@
$\frac{\partial m_{i}^{\alpha=0}(t)}{\partial\alpha}=(1-m_{i}(t)^{2})\left(-\Theta_{i}(t)+H_{i}+\sum_{j}W_{ij}m_{j}(t-1)\right),$ (162)
@@
$m_{i}(t)\approx\tanh\left(H_{i}+\sum_{j}W_{ij}m_{j}(t-1)\right).$ (164)
@@ Appendix I Auto-regression with (un)constrained least-squares
Finally, we note that in order to define the EPR, we require that the process converges to a steady-state. This requires that $\bm{-B}$ or $\bm{W-I}$ is a stable matrix i.e. all eigenvalues have negative real part. This constraint is non-convex and so cannot be enforced as part of the algorithm without resorting to more heuristic and complex optimisation methods. In this study, we found that all time-series considered, were best fit with an unconstrained model that converged to a stationary state, which did not always hold for the constrained model. As a result the EPR estimates are obtained with the unconstrained model, whilst the network inference is done with the constrained model.

### v5

@@ Abstract
The structure of a complex network plays a crucial role in determining its dynamical properties. In this work, we show that the degree to which a network is directed and hierarchically organised is closely associated with the degree to which its dynamics break detailed balance and produce entropy. We consider a range of dynamical processes and show how different directed network features affect their entropy production rate. We begin with an analytical treatment of a 2-node network followed by numerical simulations of synthetic networks using the preferential attachment and Erdös-Renyi algorithms. Next, we analyse a collection of 97 empirical networks to determine the effect of complex real-world topologies. Finally, we present a simple method for inferring broken detailed balance and directed network structure from multivariate time-series and apply our method to identify non-equilibrium dynamics and hierarchical organisation in both human neuroimaging and financial time-series. Overall, our results shed light on the consequences of directed network structure on non-equilibrium dynamics and highlight the importance and ubiquity of hierarchical organisation and non-equilibrium dynamics in real-world systems.
@@ I Introduction
As anticipated, symmetry-breaking in the organisation of complex systems is the structural feature that drives network dynamical processes out of equilibrium. Dynamical processes on directed networks differ drastically from their undirected counterparts [38, 39] including in their phase-transitions [40], synchronisation properties [41, 42], topological resilience [43, 44] and pattern formation [45, 46]. Recent studies on extensive datasets of real-world networks, have shown a ubiquity of strong directedness and clear signs of hierarchical organisation [47, 48, 39, 49, 50]. As a result, the dynamics of directed networks are more indicative of the dynamics of real-world complex systems. In particular, this strong directedness results in a marked non-normality of the operators defined on such networks [51, 48]. Consequently, non-normality can result in the underlying networked systems diverging from equilibria [48]. Motivated by the ubiquity of directed structures in the real world and their significance in non-linear dynamics, this paper aims to further strengthen this link using the formalism of stochastic thermodynamics.
@@ II Measures of directedness in networks
As anticipated, non-reciprocal interactions in complex systems disrupt detailed balance, leading to an increase in the EPR. This section delineates four established metrics to quantify the overall directedness of networks and to examine their influence on EPR within both random graphs and empirical datasets. These systems are represented as directed networks with non-negative weighted adjacency matrices, $\bm{W}=(W_{ij})$, where $W_{ij}\geq 0$ signifies the strength of the directed link $i\rightarrow j$, including self-edges $W_{ii}\geq 0$ and characterized by inherent asymmetry $\bm{W}\neq\bm{W}^{\top}$  [^1] The definition of $W_{ij}$ to represent the link from $i$ to $j$ is one of two conventions. We choose this convention to be consistent with the definition of trophic coherence [49]. The opposite, is the more typical convention for network dynamics [7].. By exploring four distinct metrics for network directedness, we seek to elucidate its effect on EPR, highlighting the emergence of hierarchical structuring within networks - an outcome deeply influenced by non-reciprocal interactions [63, 50].
@@ II.4 Parameterising directedness in networks
To investigate the dynamical effects of directedness on network dynamics, it would be preferable to continuously vary the level of asymmetry in a network. To this aim, we first generate highly hierarchical, non-normal networks using a preferential attachment (PA) scheme with weak reciprocal links (see Appendix B for details) [48]. We then linearly interpolate between this strongly directed network and its undirected Hermitian,
@@ III.1 Random walks on directed networks
which can, alternatively, be written as the Kullback-Leibler (KL) divergence between the forward and backward joint transition rates between all pairs of states,
@@ III.2 The network-based multivariate Ornstein-Uhlenbeck process
$\frac{d\bm{x}}{dt}$ $=\Theta(\gamma{\color[rgb]{0,0,0}\bm{W}^{\top}}-\bm{I})\bm{x}(t)+\bm{\nu}(t),$ (22)
@@ III.3 The Ising model
$P(\bm{x}(t+1)|\bm{x}(t))$ $=\prod_{i}\frac{\exp\left(x_{i}(t+1)\sum_{j}{\color[rgb]{0,0,0}W_{ji}}x_{j}(t)/T\right)}{2\cosh\left(\sum_{j}{\color[rgb]{0,0,0}W_{ji}}x_{j}(t)/T\right)},$ (32)
@@ III.3.2 Large networks of Ising spins
The combinatorial explosion of state space with the number of spins, means that estimating the steady state probabilities is computationally challenging for large systems. Alternative approaches include coarse-graining the state space by clustering together states and estimating transition probabilities [86, 14, 23], or using mean-field approximations [84]. Whilst using a coarse-grained state-space has been shown to capture the effect on the EPR of changing the temperature of the Ising model [23], it proved to be an inaccurate when varying the directedness of the network (see Appendix G) and so we hereby opt for a mean-field approach.
@@
$m_{i}(t)$ $\approx\tanh\sum_{j}{\color[rgb]{0,0,0}W_{ji}}m_{j}(t-1),$ (40)
$D_{il}(t)$ $\approx{\color[rgb]{0,0,0}W_{li}}(1-m^{2}_{i}(t))(1-m^{2}_{l}(t-1)).$ (41)
@@ IV.1 An exactly treatable case: 2-node networks
Finally, we can consider the Ising dynamics on this network. Whilst we do not have the exact solution, there are only four possible spin configurations in this system, therefore we can accurately estimate the steady state probabilities and consequently the EPR. The EPR, $\Phi$, for both the OU and Ising systems is plotted as a function of the network weights $(W_{12},W_{21})$ as shown in Figure 3. The positive diagonal $W_{12}=W_{21}$ represents the subset of undirected networks, which preserve detailed balance. In both cases, as the asymmetry of the network increases the EPR increases. The EPR in the OU is a function purely of the distance between the two weights, whereas for the Ising model, the relationship appears to be non-linear.
@@ IV.2 Entropy production in parameterised synthetic networks
Equipped with a range of stochastic network dynamics where we can estimate the EPR, as well as a mechanism to systematically vary the directedness of the network, we can investigate the effect of directedness on the degree of non-equilibrium. Considering networks of sizes $N=$ 10, 100, 500 and 1000, we generate hierarchical networks using preferential attachment (see Appendix B) and then interpolate between these networks and their Hermitian as described in Section II. Recall the interpolation is given by,
@@ VII Entropy production and directed networks from multivariate time-series
The key limitation of such an approach is the assumption of the linear model for potentially non-linear data. However, previous studies have found that highly non-linear systems, such as large-scale brain dynamics [108], or even chaotic systems [109], are well approximated by linear models. Furthermore, autoregressive models have previously found success in a spectrum of areas including finance and economics [57], neuroscience [110] and beyond. Finally, we note that three recent studies in neuroimaging have considered related approaches, fitting linear models to neural recordings in order to quantify the EPR [98] or the asymmetry of interactions [99, 59] but that inferring time-reversibility from incomplete linear data may yield inaccuracies [97].
@@ VII.1 Linear auto-regression of multivariate time-series
$\bm{x}(t_{i+1})$ $=[\bm{I}+\Delta t(\bm{{\color[rgb]{0,0,0}W^{\top}}-I})]\bm{x}(t_{i})+\bm{\Gamma}\bm{\zeta}_{i},$ (62)
@@ Appendix C Linear interpolation parameterises trophic directedness and irreciprocity
For the original network, $\bm{h}$ is the vector of trophic levels which solves,
@@ Appendix D Deriving the entropy production rate in the Ornstein-Uhlenbeck process
$\bm{B}$ $=\begin{pmatrix}\Theta&\Theta(1-\gamma{\color[rgb]{0,0,0}\tilde{W}_{21}}(\epsilon))\\ \Theta(1-\gamma{\color[rgb]{0,0,0}\tilde{W}_{12}}(\epsilon))&\Theta\end{pmatrix},$ (132)
@@ Appendix F Circulant networks
It is important to note that $\bm{W}$ is circulant and therefore normal. However, as it is directed, it breaks detailed balance for all the systems considered in this paper (with the exception of the RW for $N=2$). More importantly, as we will show, applying the interpolation to the cyclic networks, $\epsilon$ increases the EPR of the corresponding OU. We found in Sections V and VI that the non-normality of the underlying network was a extremely strong indicator of the EPR of the OU, yet in this case, the non-normality is 0 and the EPR still varies with $\epsilon$. This represents an important special case where hierarchical asymmetry or a global direction are not necessary for the OU to progressively break detailed balance. The lack of these cyclic structures in real-world networks [89] explains why, despite this special case, the non-normality correlates so closely with the EPR for the OU on real-world networks, as shown in Section VI.
@@
Therefore, the friction matrix $\bm{B}=\Theta(\bm{I}-\gamma\bm{{\color[rgb]{0,0,0}\hat{W}^{\top}}}(\epsilon))$ is also circulant and defined by the vector,
$\bm{b}(\epsilon)=(\Theta,-\Theta\gamma{\color[rgb]{0,0,0}\tilde{w}_{N-1}}(\epsilon),\ldots,-\Theta\gamma{\color[rgb]{0,0,0}\tilde{w}_{1}}(\epsilon)).$ (143)
@@
$\tilde{b}_{k}$ $=\Theta-\Theta\gamma\sum_{j=1}^{N-1}{\color[rgb]{0,0,0}\tilde{w}_{N-j}}(\epsilon)\exp{\bm{i}\frac{2\pi jk}{N}}.$ (145)
@@
$\tilde{b}_{k}$ $=\Theta-\Theta\gamma\sum_{j=1}^{N-1}(\epsilon{\color[rgb]{0,0,0}w_{N-j}})(\bm{i}\sin(\frac{2\pi jk}{N})+\cos(\frac{2\pi jk}{N}))-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(1-\epsilon)(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N}),$ (148)
@@
$\Phi(\epsilon)$ $=\sum_{k=0}^{N-1}\frac{\Theta^{2}\gamma^{2}\left(\sum_{j=1}^{N-1}\epsilon{\color[rgb]{0,0,0}w_{N-j}}\sin(\frac{2\pi jk}{N})\right)^{2}}{\Theta-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(1-\epsilon)(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N})-\Theta\gamma\sum_{j=1}^{N-1}\epsilon{\color[rgb]{0,0,0}w_{N-j}}\cos(\frac{2\pi jk}{N})},$ (149)
$=\epsilon^{2}\sum_{k=0}^{N-1}\frac{\Theta^{2}\gamma^{2}\left(\sum_{j=1}^{N-1}{\color[rgb]{0,0,0}w_{N-j}}\sin(\frac{2\pi jk}{N})\right)^{2}}{\Theta-\Theta\gamma\sum_{j=1}^{\lfloor\frac{N}{2}\rfloor}(w_{j}+w_{N-j})\cos(\frac{2\pi jk}{N})},$ (150)
@@ Appendix H Derivation of the naive mean field for the asymmetric Ising model
$h^{\alpha}_{i}(t+1)=(1-$ $\alpha)\Theta_{i}(t)+\alpha\left(\frac{1}{T}(H_{i}+\sum_{j}{\color[rgb]{0,0,0}W_{ji}}x_{j}(t))\right).$ (159)
@@
$\frac{\partial m_{i}^{\alpha=0}(t)}{\partial\alpha}=(1-m_{i}(t)^{2})\left(-\Theta_{i}(t)+H_{i}+\sum_{j}{\color[rgb]{0,0,0}W_{ji}}m_{j}(t-1)\right),$ (162)
@@
$m_{i}(t)\approx\tanh\left(H_{i}+\sum_{j}{\color[rgb]{0,0,0}W_{ji}}m_{j}(t-1)\right).$ (164)
@@ Appendix I Auto-regression with (un)constrained least-squares
Finally, we note that in order to define the EPR, we require that the process converges to a steady-state. This requires that $\bm{-B}$ or $\bm{{\color[rgb]{0,0,0}W^{\top}}-I}$ is a stable matrix i.e. all eigenvalues have negative real part. This constraint is non-convex and so cannot be enforced as part of the algorithm without resorting to more heuristic and complex optimisation methods. In this study, we found that all time-series considered, were best fit with an unconstrained model that converged to a stationary state, which did not always hold for the constrained model. As a result the EPR estimates are obtained with the unconstrained model, whilst the network inference is done with the constrained model.